OpenStudy (anonymous):

How would you find the surface area of a sphere if the circumference is 3pi

OpenStudy (anonymous):

the surface area is 4*pi*radius^2, the circumference is 2*pi*radius, which in your case is 3pi so radius is 3/2 and then square the radius and multiply by 4 and pi :)

OpenStudy (anonymous):

Find the radius r and plug the result into the expression for the sphere's surface area. \[4*\pi*r^2\] The ratio of the circumfrence to the diameter is equal to Pi. The diameter is twice the radius. \[{3 \pi \over 2r} = \pi\] r is 3/2 so the surface area is \[4\pi({3\over2})^2 = 9\pi\]